Hi
A V and W had a total of 510 stickers. The ratio of A to V stickers was 6:7 at first. After A and V each gave away half of their stickers the 3 had 345 stickers left. How many did W have at first.
Thank you

Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

A + V + W = 510

510 - 1/2 * (A + V) = 345

this means that 1/2 * (A + V) = 165, because 510 - 165 = 345.

this means that (A + V) = 330

since A + V + W = 510, then:

W must be equal to 510 - 330 = 180.

that should be your answer.

the additional information that A/V = 6/7 looks like it was extraneous, as it does not appear that it had anything to do with solving the problem.





Answer by ikleyn(52818)

.
A V and W had a total of 510 stickers. The ratio of A to V stickers was 6:7 at first. After A and V each
gave away half of their stickers the 3 had 345 stickers left. How many did W have at first.
~~~~~~~~~~~~~~~~~


                As a continuation of the solution by @Theo.
            ===============================================



In this problem,  the imposed conditions  CONTRADICT  each to the other,  so such integer numbers,
satisfying the imposed conditions,  DO  NOT  EXIST.


Indeed,  if the sum   A + V = 330,   as  @Theo correctly found in his solution,  then the ratio of  A  to  V  stickers  CAN  NOT  be   6:7,

since there are no such integers  A  and  V with the sum of  330  and with the ratio  6:7.


Thus this problem is a  FAKE,  good only for re-cycling.